OpenStudy (hkmiu):

Identify the type of function that models the labor cost for repairing a computer if the charge is $50 per hour or fraction thereof. a.greatest integer function c.absolute value function b.step function

OpenStudy (anonymous):

it is a step function for sure

OpenStudy (anonymous):

the greatest integer function is also a step function, but not this step function

OpenStudy (hkmiu):

That gets me confused :O ..

OpenStudy (anonymous):

then ignore my last remarks this function looks like \(50\) for \(t<1\) then \(100\) for \(1<t\leq 2\) etc in other words a step function

OpenStudy (hkmiu):

The cost of mailing a letter is $0.37 for the first ounce and $0.23 for each additional ounce or portion thereof. then this would be a greatest integer function right?

OpenStudy (anonymous):

no

OpenStudy (hkmiu):

@________@

OpenStudy (anonymous):

greatest integer function is \(0\) for \(0<x\leq 1\) then \(1\) for \(1\leq x<2\) and so on

OpenStudy (anonymous):

the greatest integer function is a kind of step function, it is a specific step function not just any step function

OpenStudy (hkmiu):

I'll look back onto that. Thank you. I'm probably not connecting everything together bc im so tired lol. I did math for 8 hours straight.

OpenStudy (anonymous):

take a break

OpenStudy (hkmiu):

Dinner is soon :) so yes.